# MaxSup: Overcoming Representation Collapse in Label Smoothing

### v2

@@ Abstract
Label Smoothing (LS) is widely adopted to reduce overconfidence in neural network predictions and improve generalization. Despite these benefits, recent studies reveal two critical issues with LS. First, LS induces overconfidence in misclassified samples. Second, it compacts feature representations into overly tight clusters, diluting intra-class diversity, although the precise cause of this phenomenon remained elusive. In this paper, we analytically decompose the LS-induced loss, exposing two key terms: (i) a regularization term that dampens overconfidence only when the prediction is correct, and (ii) an error-amplification term that arises under misclassifications. This latter term compels the network to reinforce incorrect predictions with undue certainty, exacerbating representation collapse. To address these shortcomings, we propose Max Suppression (MaxSup), which applies uniform regularization to both correct and incorrect predictions by penalizing the top-1 logit rather than the ground-truth logit. Through extensive feature-space analyses, we show that MaxSup restores intra-class variation and sharpens inter-class boundaries. Experiments on large-scale image classification and multiple downstream tasks confirm that MaxSup is a more robust alternative to LS, consistently reducing overconfidence while preserving richer feature representations.  [^1] Code is available at: https://anonymous.4open.science/r/Maximum-Suppression-Regularization-DB0C.
@@ 1 Introduction
Multi-class classification [17, 24] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal. In practice, however, classes often share low-level features [29, 40] or exhibit high-level semantic similarities [3, 22, 38], rendering the one-hot assumption overly restrictive. Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [10].
To moderate overconfidence, Szegedy et al. [32] introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class. LS has since become prevalent in image recognition [11, 20, 33, 43] and neural machine translation [1, 6], often boosting accuracy and calibration [21]. Yet subsequent work indicates that LS can overly compress features into tight clusters [13, 26, 37], hindering intra-class variability and transferability [5]. In parallel, Zhu et al. [44] found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
In this paper, we reveal that LS’s training objective inherently contains an error amplification term. This term pushes the network to reinforce incorrect predictions with exaggerated certainty, yielding highly confident misclassifications and further compressing feature clusters (Section 3.1, Table 1). Building on Zhu et al. [44], we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics. Through our analysis, we further show that punishing the ground-truth logit during misclassification reduces intra-class variation (Table 2), a phenomenon corroborated by Grad-CAM visualizations (Figure 2).
To overcome these shortcomings, we introduce Max Suppression (MaxSup), a method that retains the beneficial regularization effect of LS while eliminating its error amplification. Rather than penalizing the ground-truth logit, MaxSup focuses on the model’s top-1 logit, ensuring a consistent regularization signal regardless of whether the current prediction is correct or misclassified. By preserving the ground-truth logit in misclassifications, MaxSup sustains richer intra-class variability and sharpens inter-class boundaries. As visualized in Figure 1, this approach mitigates the feature collapse and attention drift often induced by LS, ultimately leading to more robust representations. Through comprehensive experiments in both image classification (Section 4.2) and semantic segmentation (Section 4.3), we show that MaxSup not only alleviates severe intra-class collapse but also consistently boosts top-1 accuracy (+0.4%–1.0% improvement over LS on ImageNet) and robustly enhances downstream transfer performance (Section 4.1).
@@
  • We demonstrate superior performance across tasks and architectures, including ResNet, MobileNetV2, and DeiT-S, where MaxSup significantly boosts accuracy on ImageNet (up to +1.0% in top-1) and consistently delivers stronger representations for downstream tasks such as semantic segmentation and robust transfer learning.
@@ 2.1 Regularization
Regularization techniques aim to improve the generalization of deep neural networks by constraining model complexity. Classical methods like $\ell_{2}$ [16] and $\ell_{1}$ [45] impose direct penalties on large or sparse weights, while Dropout [30] randomly deactivates neurons to discourage over-adaptation. In the realm of loss-based strategies, Label Smoothing (LS) [32] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [21]. Variants such as Online Label Smoothing (OLS) [41] and Zipf Label Smoothing (Zipf-LS) [19] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions. However, they do not fully address the fundamental issue that emerges when the ground-truth logit is not the highest one (see Section 3.1, Table 1). Other loss-based regularizers focus on alternative aspects of the predictive distribution. Confidence Penalty [23] penalizes the model’s confidence directly, while Logit Penalty [4] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [13]. Despite these benefits, Logit Penalty can inadvertently shrink intra-class variation, thereby hampering transfer learning (see Section 4.1). Unlike the aforementioned methods, MaxSup enforces regularization by penalizing only the top-1 logit $z_{\textit{max}}$ rather than the ground-truth logit $z_{gt}$. In LS-based approaches, suppressing $z_{gt}$ for misclassified samples can worsen errors, whereas MaxSup applies a uniform penalty regardless of whether the model’s prediction is correct. Consequently, MaxSup avoids the error amplification effect, retains richer intra-class diversity (see Table 2), and achieves robust transfer performance across diverse datasets and model families (see Table 3).
@@ 2.2 Studies on Label Smoothing
Label Smoothing has also been studied extensively under knowledge distillation. For instance, Yuan et al. [39] observed that LS can approximate the effect of a teacher–student framework, while Shen et al. [28] investigated its role in such pipelines more systematically. Additionally, Chandrasegaran et al. [2] demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes. Concurrently, Kornblith et al. [13] showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance. From a Neural Collapse perspective [42, 9], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in Xu and Liu [37]. Our goal is to overcome LS’s inherent error amplification effect. Rather than adjusting how the smoothed label distribution is constructed (as in OLS or Zipf-LS), MaxSup directly penalizes the highest logit $z_{\textit{max}}$. This design ensures consistent regularization even if $z_{gt}$ is not the top logit, thereby avoiding the degradation in performance typical of misclassified samples under LS (see Section 3.2). Moreover, MaxSup integrates seamlessly into standard training pipelines, introducing negligible computational overhead beyond substituting the LS term.
@@ 3 Max Suppression Regularization (MaxSup)
We first partition the training objective into two components: the standard Cross-Entropy (CE) loss and a regularization term introduced by Label Smoothing (LS). By expressing LS in terms of logits (Theorem 3.3), we isolate two key factors: a regularization term that controls overconfidence and an error amplification term that enlarges the gap between the ground-truth logit $z_{gt}$ and any higher logits (Corollary 3.4, Equation 5), ultimately degrading performance. To address these issues, we propose Max Suppression Regularization (MaxSup), which applies the penalty to the largest logit $z_{\textit{max}}$ rather than $z_{gt}$ (Equation 8, Section 3.2). This shift delivers consistent regularization for both correct and incorrect predictions, preserves intra-class variation, and bolsters inter-class separability. Consequently, MaxSup mitigates the representation collapse found in LS, attains superior ImageNet-1K accuracy (Table 1), and improves transferability (Table 2, Table 3). The following sections elaborate on MaxSup’s formulation and integration into the training pipeline.
@@ 3.1 Revisiting Label Smoothing
Label Smoothing (LS) is a regularization technique designed to reduce overconfidence by softening the target distribution. Rather than assigning probability $1$ to the ground-truth class and 00 to all others, LS redistributes a fraction $\alpha$ of the probability uniformly across all classes:
@@
Where, $\mathbf{q}$ is the predicted probability vector, $H(\cdot)$ denotes the Cross-Entropy, and $\frac{\mathbf{1}}{K}$ is the uniform distribution introduced by LS. This shows that LS adds a regularization term, $L_{\textit{LS}}$, which smooths the output distribution and helps to reduce overfitting. (See Appendix A for a formal proof.)
Building on Lemma 3.2, we next explicitly express $L_{\textit{LS}}$ at the logit level for further analysis.
@@
where $z_{gt}$ is the logit corresponding to the ground-truth class, and $\tfrac{1}{K}\sum_{k=1}^{K}z_{k}$ is the average logit. Thus, LS penalizes the gap between $z_{gt}$ and the average logit, encouraging a more balanced output distribution and reducing overconfidence. (See Appendix B for the proof.)
@@
where $M$ and $N$ are the numbers of logits below and above $z_{gt}$, respectively ($M+N=K-1$). Note that the error amplification term vanishes when $z_{gt}=z_{\textit{max}}$.
@@
These issues with LS on misclassified samples have also been systematically observed in prior work [35]. By precisely disentangling these two components (regularization vs. error amplification), we can design a more targeted and effective solution.
@@
Ablation Study on LS Components. To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [33] without Mixup or CutMix. As indicated in Table 1, the performance gains from LS stem solely from the regularization term, whereas the error amplification term degrades accuracy. In contrast, our MaxSup omits the error amplification component and leverages only the beneficial regularization, thereby boosting accuracy beyond that of standard LS. Specifically, Table 1 shows that LS’s overall improvement can be attributed exclusively to its regularization contribution; the error amplification term consistently reduces accuracy (e.g., to 73.63% or 73.69%). Disabling only the error amplification while retaining the regularization yields a slight but measurable improvement (75.98% vs. 75.91%). By fully removing error amplification and faithfully preserving the helpful aspects of LS, our MaxSup achieves 76.12% accuracy, clearly and consistently outperforming LS. This result underscores that MaxSup directly tackles LS’s fundamental shortcoming by maintaining a consistent and meaningful regularization signal—even when the top-1 prediction is incorrect.
@@ 3.2 Max Suppression Regularization
Building on our analysis in Section 3.1, we find that Label Smoothing (LS) not only impacts correctly classified samples but also influences misclassifications in unintended and harmful ways. Specifically, LS suffers from two main limitations: inconsistent regularization and error amplification. As illustrated in Table 1, LS penalizes the ground-truth logit $z_{gt}$ even in misclassified examples, needlessly widening the gap between $z_{gt}$ and the erroneous top-1 logit. To resolve these critical shortcomings, we propose Max Suppression Regularization (MaxSup), which explicitly penalizes the largest logit $z_{\textit{max}}$ rather than $z_{gt}$. This key design choice ensures uniform regularization across both correct and misclassified samples, effectively eliminating the error-amplification issue in LS (Table 1) and preserving the ground-truth logit’s integrity for more stable, robust learning.
@@
$\underbrace{H(\mathbf{s},\mathbf{q})}_{\text{CE with Soft Labels}}\;=\;\underbrace{H(\mathbf{y},\mathbf{q})}_{\text{CE with Hard Labels}}\;+\;\underbrace{L_{\textit{MaxSup}}}_{\text{Max Suppression Loss}},$ (6)
@@
Logit-Level Formulation of MaxSup. Building on the logit-level perspective introduced for LS in Section 3.1, we can express $L_{\textit{MaxSup}}$ as:
@@
Comparison with Label Smoothing. MaxSup fundamentally differs from LS in handling correct and incorrect predictions. When $z_{gt}=z_{\textit{max}}$, both LS and MaxSup similarly reduce overconfidence. However, when $z_{gt}\neq z_{\textit{max}}$, LS shrinks $z_{gt}$, widening the gap with the incorrect logit, whereas MaxSup penalizes $z_{\textit{max}}$, preserving $z_{gt}$ from undue suppression. As illustrated in Figure 2, this helps the model recover from mistakes more effectively and avoid reinforcing incorrect predictions.
@@
Thus, the top-1 logit $z_{\textit{max}}$ is reduced by $\alpha\bigl{(}1-\tfrac{1}{K}\bigr{)}$, while all other logits slightly increase by $\tfrac{\alpha}{K}$. In misclassified cases, the ground-truth logit $z_{gt}$ is spared from penalization, avoiding the error-amplification issue seen in LS. For completeness, Appendix A provides the full gradient derivation.

### v3

@@ Abstract
Label Smoothing (LS) is widely adopted to reduce overconfidence in neural network predictions and improve generalization. Despite these benefits, recent studies reveal two critical issues with LS. First, LS induces overconfidence in misclassified samples. Second, it compacts feature representations into overly tight clusters, diluting intra-class diversity, although the precise cause of this phenomenon remained elusive. In this paper, we analytically decompose the LS-induced loss, exposing two key terms: (i) a regularization term that dampens overconfidence only when the prediction is correct, and (ii) an error-amplification term that arises under misclassifications. This latter term compels the network to reinforce incorrect predictions with undue certainty, exacerbating representation collapse. To address these shortcomings, we propose Max Suppression (MaxSup), which applies uniform regularization to both correct and incorrect predictions by penalizing the top-1 logit rather than the ground-truth logit. Through extensive feature-space analyses, we show that MaxSup restores intra-class variation and sharpens inter-class boundaries. Experiments on large-scale image classification and multiple downstream tasks confirm that MaxSup is a more robust alternative to LS.  [^1] https://github.com/ZhouYuxuanYX/Maximum-Suppression-Regularization.
@@ 1 Introduction
Multi-class classification [lecun1998mnist, russakovsky2015imagenet] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal. In practice, however, classes often share low-level features [silla2011survey, zeiler2014visualizing] or exhibit high-level semantic similarities [chen2021hsva, novack2023chils, yi2022exploring], rendering the one-hot assumption overly restrictive. Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [guo2020online].
To moderate overconfidence, szegedy2016rethinking introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class. LS has since become prevalent in image recognition [he2016deep, liu2021swin, touvron2021training, zhou2022sp] and neural machine translation [alves2023steering, gao2020towards], often boosting accuracy and calibration [muller2019does]. Yet subsequent work indicates that LS can overly compress features into tight clusters [kornblith2021better, sariyildiz2022no, xu2023quantifying], hindering intra-class variability and transferability [feng2021rethinking]. In parallel, zhu2022rethinking found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
In this paper, we reveal that LS’s training objective inherently contains an error amplification term. This term pushes the network to reinforce incorrect predictions with exaggerated certainty, yielding highly confident misclassifications and further compressing feature clusters (Section˜3.1, Table˜1). Building on zhu2022rethinking, we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics. Through our analysis, we further show that punishing the ground-truth logit during misclassification reduces intra-class variation (Table˜2), a phenomenon corroborated by Grad-CAM visualizations (Figure˜2).
To overcome these shortcomings, we introduce Max Suppression (MaxSup), a method that retains the beneficial regularization effect of LS while eliminating its error amplification. Rather than penalizing the ground-truth logit, MaxSup focuses on the model’s top-1 logit, ensuring a consistent regularization signal regardless of whether the current prediction is correct or misclassified. By preserving the ground-truth logit in misclassifications, MaxSup sustains richer intra-class variability and sharpens inter-class boundaries. As visualized in Figure˜1, this approach mitigates the feature collapse and attention drift often induced by LS, ultimately leading to more robust representations. Through comprehensive experiments in both image classification (Section 4.2) and semantic segmentation (Section 4.3), we show that MaxSup not only alleviates severe intra-class collapse but also consistently boosts top-1 accuracy and robustly enhances downstream transfer performance (Section 4.1).
@@
  • We demonstrate superior performance across tasks and architectures, including ResNet, MobileNetV2, and DeiT-S, where MaxSup significantly boosts accuracy on ImageNet and consistently delivers stronger representations for downstream tasks such as semantic segmentation and robust transfer learning.
@@ 2.1 Regularization
Regularization techniques aim to improve the generalization of deep neural networks by constraining model complexity. Classical methods like $\ell_{2}$ [krogh1991simple] and $\ell_{1}$ [zou2005regularization] impose direct penalties on large or sparse weights, while Dropout [Srivastava2014] randomly deactivates neurons to discourage over-adaptation. In the realm of loss-based strategies, Label Smoothing (LS) [szegedy2016rethinking] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [muller2019does]. Variants such as Online Label Smoothing (OLS) [zhang2021delving] and Zipf Label Smoothing (Zipf-LS) [liang2022efficient] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions. However, they do not fully address the fundamental issue that emerges when the ground-truth logit is not the highest one (see Section˜3.1, Table 1). Other loss-based regularizers focus on alternative aspects of the predictive distribution. Confidence Penalty [pereyra2017regularizing] penalizes the model’s confidence directly, while Logit Penalty [dauphin2021deconstructing] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [kornblith2021better]. Despite these benefits, Logit Penalty can inadvertently shrink intra-class variation, thereby hampering transfer learning (see Section˜4.1). Unlike the aforementioned methods, MaxSup enforces regularization by penalizing only the top-1 logit $z_{\textit{max}}$ rather than the ground-truth logit $z_{gt}$. In LS-based approaches, suppressing $z_{gt}$ for misclassified samples can worsen errors, whereas MaxSup applies a uniform penalty regardless of whether the model’s prediction is correct. Consequently, MaxSup avoids the error amplification effect, retains richer intra-class diversity (see Table 2), and achieves robust transfer performance across diverse datasets and model families (see Table 3).
@@ 2.2 Studies on Label Smoothing
Label Smoothing has also been studied extensively under knowledge distillation. For instance, yuan2020revisiting observed that LS can approximate the effect of a teacher–student framework, while shen2021label investigated its role in such pipelines more systematically. Additionally, chandrasegaran2022revisiting demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes. Concurrently, kornblith2021better showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance. From a Neural Collapse perspective [zhou2022all, guo2024cross], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in xu2023quantifying. Our goal is to overcome LS’s inherent error amplification effect. Rather than adjusting how the smoothed label distribution is constructed (as in OLS or Zipf-LS), MaxSup directly penalizes the highest logit $z_{\textit{max}}$. This design ensures consistent regularization even if $z_{gt}$ is not the top logit, thereby avoiding the degradation in performance typical of misclassified samples under LS (see Section˜3.2). Moreover, MaxSup integrates seamlessly into standard training pipelines, introducing negligible computational overhead beyond substituting the LS term.
@@ 3 Max Suppression Regularization (MaxSup)
We first partition the training objective into two components: the standard Cross-Entropy (CE) loss and a regularization term introduced by Label Smoothing (LS). By expressing LS in terms of logits (Theorem˜3.3), we isolate two key factors: a regularization term that controls overconfidence and an error amplification term that enlarges the gap between the ground-truth logit $z_{gt}$ and any higher logits (Corollary˜3.4, Equation˜5), ultimately degrading performance. To address these issues, we propose Max Suppression Regularization (MaxSup), which applies the penalty to the largest logit $z_{\textit{max}}$ rather than $z_{gt}$ (Equation˜8, Section˜3.2). This shift delivers consistent regularization for both correct and incorrect predictions, preserves intra-class variation, and bolsters inter-class separability. Consequently, MaxSup mitigates the representation collapse found in LS, attains superior ImageNet-1K accuracy (Table˜1), and improves transferability (Table˜2, Table˜3). The following sections elaborate on MaxSup’s formulation and integration into the training pipeline.
@@ 3.1 Revisiting Label Smoothing
Label Smoothing (LS) is a regularization technique designed to reduce overconfidence by softening the target distribution. Rather than assigning probability $1$ to the ground-truth class and 0 to all others, LS redistributes a fraction $\alpha$ of the probability uniformly across all classes:
@@
Where, $\mathbf{q}$ is the predicted probability vector, $H(\cdot)$ denotes the Cross-Entropy, and $\frac{\mathbf{1}}{K}$ is the uniform distribution introduced by LS. This shows that LS adds a regularization term, $L_{\textit{LS}}$, which smooths the output distribution and helps to reduce overfitting. (See Appendix˜B for a formal proof.)
Building on Lemma˜3.2, we next explicitly express $L_{\textit{LS}}$ at the logit level for further analysis.
@@
where $z_{gt}$ is the logit corresponding to the ground-truth class, and $\tfrac{1}{K}\sum_{k=1}^{K}z_{k}$ is the average logit. Thus, LS penalizes the gap between $z_{gt}$ and the average logit, encouraging a more balanced output distribution and reducing overconfidence. (See Appendix˜C for the proof.)
@@
where $M$ and $N$ are the numbers of logits below and above $z_{gt}$, respectively (M+N=K1M+N=K-1). Note that the error amplification term vanishes when $z_{gt}=z_{\textit{max}}$.
@@
These issues with LS on misclassified samples have also been systematically observed in prior work [xia2024understanding]. By precisely disentangling these two components (regularization vs. error amplification), we can design a more targeted and effective solution.
@@
Ablation Study on LS Components. To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [touvron2021training] without Mixup or CutMix. As indicated in Table˜1, the performance gains from LS stem solely from the regularization term, whereas the error amplification term degrades accuracy. In contrast, our MaxSup omits the error amplification component and leverages only the beneficial regularization, thereby boosting accuracy beyond that of standard LS. Specifically, Table˜1 shows that LS’s overall improvement can be attributed exclusively to its regularization contribution; the error amplification term consistently reduces accuracy (e.g., to 73.63% or 73.69%). Disabling only the error amplification while retaining the regularization yields a slight but measurable improvement (75.98% vs. 75.91%). By fully removing error amplification and faithfully preserving the helpful aspects of LS, our MaxSup achieves 76.12% accuracy, clearly and consistently outperforming LS. This result underscores that MaxSup directly tackles LS’s fundamental shortcoming by maintaining a consistent and meaningful regularization signal—even when the top-1 prediction is incorrect.
@@ 3.2 Max Suppression Regularization
Building on our analysis in Section˜3.1, we find that Label Smoothing (LS) not only impacts correctly classified samples but also influences misclassifications in unintended and harmful ways. Specifically, LS suffers from two main limitations: inconsistent regularization and error amplification. As illustrated in Table 1, LS penalizes the ground-truth logit $z_{gt}$ even in misclassified examples, needlessly widening the gap between $z_{gt}$ and the erroneous top-1 logit. To resolve these critical shortcomings, we propose Max Suppression Regularization (MaxSup), which explicitly penalizes the largest logit $z_{\textit{max}}$ rather than $z_{gt}$. This key design choice ensures uniform regularization across both correct and misclassified samples, effectively eliminating the error-amplification issue in LS (Table 1) and preserving the ground-truth logit’s integrity for more stable, robust learning.
@@
[…]
@@
Logit-Level Formulation of MaxSup. Building on the logit-level perspective introduced for LS in Section˜3.1, we can express $L_{\textit{MaxSup}}$ as:
@@
Comparison with Label Smoothing. MaxSup fundamentally differs from LS in handling correct and incorrect predictions. When $z_{gt}=z_{\textit{max}}$, both LS and MaxSup similarly reduce overconfidence. However, when $z_{gt}\neq z_{\textit{max}}$, LS shrinks $z_{gt}$, widening the gap with the incorrect logit, whereas MaxSup penalizes $z_{\textit{max}}$, preserving $z_{gt}$ from undue suppression. As illustrated in Figure˜2, this helps the model recover from mistakes more effectively and avoid reinforcing incorrect predictions.
@@
Thus, the top-1 logit $z_{\textit{max}}$ is reduced by $\alpha\bigl(1-\tfrac{1}{K}\bigr)$, while all other logits slightly increase by $\tfrac{\alpha}{K}$. In misclassified cases, the ground-truth logit $z_{gt}$ is spared from penalization, avoiding the error-amplification issue seen in LS. For completeness, Appendix B provides the full gradient derivation. While [xia2024understanding] conducted a related gradient analysis of the training loss, it focuses specifically on the setting of selective classification, and examines a posthoc logit normalization technique to mitigate confidence calibration issues. However, this approach addresses only the overconfidence problem of label smoothing (LS), without tackling representation collapse. Moreover, our work presents a logit-level reformulation of LS that provides a deeper theoretical understanding of why LS amplifies errors.